Assume C-50+0.6 (Y -T) G = 15 I = 15 T=2 (a) Y at equilibrium is equal to (b) C at equilibrium is equal to
Answer #1

Answer

Aggregate Expenditure(AE) is given by:

AE = C + I + G. Here C = 50 + 0.6(Y - T) , I = 15 , G = 15 and T = 2

So, AE = 50 + 0.6(Y - 2) + 15 + 15 = 78.8 + 0.6Y

Equilibrium occurs when Income(Y) = Aggregate Expenditure

Hence at equilibrium, Y = AE

=> Y =  78.8 + 0.6Y

=> Y = 197

Y at equilibrium is equal to 197

(b) At equilibrium, C =  50 + 0.6(Y - T) = 50 + 0.6(197 - 2) = 167

Hence C at equilibrium is equal to 167